OpenStudy (anonymous):

Find the anti-derivative of xsqrtx+1 using substitution show steps

hartnn (hartnn):

are u sure its \(f(x)=x\sqrt x+1\) ?

OpenStudy (anonymous):

\[x \sqrt{x+1}\]

hartnn (hartnn):

ok, then put x+1 = u find du=.. ?

OpenStudy (anonymous):

du=((x^2)/2)+x ?

hartnn (hartnn):

x+1 = u taking deribative dx+0 = du so, dx is just du.

OpenStudy (anonymous):

oooo I was doing the anti derivative for du. ok thanks

hartnn (hartnn):

hmm....so could you solove further ?

OpenStudy (anonymous):

so now the equation is sqrtu du ?

hartnn (hartnn):

\(\sqrt{x+1} = \sqrt u \\ x=u-1 \\dx=du\) so whats your new integral ?

OpenStudy (anonymous):

u-1sqrtu du

hartnn (hartnn):

yes, now thats an easy integral t solve can u ?

OpenStudy (anonymous):

yea I got this thanks

hartnn (hartnn):

welcome ^_^